use Leibnitz theorem to evaluate the fourth derivative of \\(\\left(2x^{3}+x^{2}+x+2\\right)e^{2x}\\)
1
Expert's answer
2019-10-21T08:32:45-0400

Lebnicz formula for the product


"(f(x)g(x)) ^{(n)}= \sum\limits _{r=0}^{n}C^r_n(f )^{(n-r)}(x) (g )^{(r)}(x)"


"f(x)= 2x^{3}+x^{2}+x+2, g(x)=e^{2x}"


"(f(x)g(x)) ^{(4)}=((2x^{3}+x^{2}+x+2)e^{2x}) ^{(4)} = \\ 
 = \sum \limits _{r=0}^{4} C^r_4(2x^{3}+x^{2}+x+2) ^{(4-r)} (e^{2x} )^{(r)}" =

"=\sum\limits _{r=0}^{n}C^r_4((2x^{3}+x^{2}+x+2) )^{(4-r)}(x) (e^{2x} )^{(r)}=\\=C^0_4((2x^{3}+x^{2}+x+2) )^{(4)}(x) (e^{2x} )+\\
+C^1_4((2x^{3}+x^{2}+x+2) )^{(3)}(x) (e^{2x} )^{(1)}+\\+C^2_4((2x^{3}+x^{2}+x+2) )^{(2)}(x) (e^{2x} )^{(2)}+\\+C^3_4((2x^{3}+x^{2}+x+2) )^{(1)}(x) (e^{2x} )^{(3)}+\\+C^4_4((2x^{3}+x^{2}+x+2) )(x) (e^{2x} )^{(4)}"


"\frac{}{}"   "C^r_n=\\frac{n!}{(n-k)! k!} \n\u200b"

"C^0_4=\frac{4!}{(4)! 0!}=1 \\ C^1_4=\frac{4!}{(3)! 1!}=4\\ C^2_4=\frac{4!}{(2)! 2!}=6\\C^3_4=\frac{4!}{(1)! 3!}=4\\C^4_4=\frac{4!}{(0)! 4!}=1"

"(2x^{3}+x^{2}+x+2) ^{(1)}=6x^{2}+2x^+1\\
(2x^{3}+x^{2}+x+2) ^{(2)}=(6x^{2}+2x+1)^{(1)}=12x+2\\
(2x^{3}+x^{2}+x+2) ^{(3)}=(12x+2)^{(1)}=12\\
(2x^{3}+x^{2}+x+2) ^{(4)}=0"


"(e^{2x} )^{(1)}=2e^{2x} \\
(e^{2x} )^{(2)}=4e^{2x}\\
(e^{2x} )^{(3)}=8e^{2x}\\
(e^{2x} )^{(4)}=16e^{2x}"

So


"((2x^{3}+x^{2}+x+2)e^{2x}) ^{(4)}=\\=1*0*e^{2x}+4*2*12 e^{2x}+ \\+6*4*(12x+2) e^{2x}+ \\+4*8*(6x^{2}+2x+1) e^{2x}+\\+1*16 (2x^{3}+x^{2}+x+2) e^{2x}=\\=e^{2x}(32x^3+208x^2+368x+208)"


Answer:


"\frac{d^4}{dx^4} ((2x^{3}+x^{2}+x+2)e^{2x})=e^{2x}(32x^3+208x^2+368x+208)."